OpenStudy (anonymous):

find the slope intercept equation of the line that has the given characteristics. slope 0 and y-intercept (0,9)

OpenStudy (anonymous):

@Abmon98

OpenStudy (abmon98):

y-9=0(x-0) y-9=0 y=9

OpenStudy (anonymous):

i have to write in y=mx+b form.

OpenStudy (abmon98):

y=0(0)+9

OpenStudy (texaschic101):

if the slope is 0 then it is a horizontal line. That means y = 9 OR y = 0(x) + 9